The length of a rectangle is 5 ft less than double the width, and the area of the rectangle is 33 ft 2. Find the dimensions of the rectangle.
Answer by TimothyLamb(4379)

Lw = 32
L = 2w - 5
---
Lw = 32
(2w - 5)w = 32
2ww - 5w - 32 = 0
---
the above quadratic equation is in standard form, with a=2, b=-5 and c=-32
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 -5 -32
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
w = 5.44076365
w = -2.94076365
---
the negative root doesn't fit the problem statement, so use the positive root:
---
w = 5.44076365
L = 2w - 5
L = 2*5.44076365 - 5
L = 5.8815273
---
answer:
w = 5.44076365 ft
L = 5.8815273 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php